OpenStudy (anonymous):

What is the solution to the equation X/9 = 3? x = 1 x = 3 x = 9 x = 27

OpenStudy (anonymous):

= 3??

OpenStudy (anonymous):

It's 3

OpenStudy (mertsj):

Multiply both sides by 9 and see what happens

OpenStudy (anonymous):

9/3 = 3

OpenStudy (mertsj):

You mean it's 3 because 3 divided by 9 is 3?

OpenStudy (mertsj):

x/9 means x divided by 9 not 9 divided by x

OpenStudy (anonymous):

ok well sorry 3/9 = 3

OpenStudy (anonymous):

same diff

OpenStudy (anonymous):

27

OpenStudy (anonymous):

fail

OpenStudy (mertsj):

3/9=1/3 NOT 3

OpenStudy (anonymous):

this question is hurting my brain

OpenStudy (mertsj):

If you have 3 dollars and have to divide them equally among 9 people they won't each get 3 dollars.

OpenStudy (anonymous):

anuragtripathi is right

OpenStudy (anonymous):